OpenStudy (anonymous):

What is the standard form of 3−i2i?

OpenStudy (anonymous):

3-i over 2i like a fraction

OpenStudy (itsbribro):

\[\frac{ 3-i }{ 2i }\]

OpenStudy (itsbribro):

like this?

OpenStudy (trojanpoem):

standard form is a + bi conjunct of a + bi = a - bi ( not -a + bi) to reach that, multiply times the conjunct of the denominator: 3- i /2i -2i(3-i) / -4i^2 i = sqrt(-1) i^2 = -1 -6i -2/ 4 -3/2i - 0.5